OpenStudy (anonymous):

Determine whether this series converges or diverges. If it converges find the sum (see contents of quest for series)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ 1 }{ 2n(n+1)}\]

OpenStudy (abb0t):

Use one of the tests. When in doubt, use ratio test. Always works.

OpenStudy (abb0t):

Usually works*

OpenStudy (anonymous):

i am not sure of my working. i first split it into partial fractions then took the limit to infinity is that correct?

OpenStudy (abb0t):

According to ratio test: \[L = \lim_{n \rightarrow ∞}\left| \frac{ a_{n+1} }{ a_n } \right|\] Converges (ans converges absolutely) if L < 1 Diverges if L > 1 Use different test if L = 1

OpenStudy (zzr0ck3r):

Partial fractions, then it looks as if its telescopic.

OpenStudy (anonymous):

partial fractions give me \[\frac{ 1 }{ 2n }-\frac{ 0.5 }{ n+1 }\]

OpenStudy (zzr0ck3r):

Yes thats it its telescopic. U understand?

OpenStudy (anonymous):

no, explain further

OpenStudy (anonymous):

AST works on series that alternates between positive and neagtive. is the above series alternating

OpenStudy (zzr0ck3r):

Converges to 1/2

OpenStudy (zzr0ck3r):

Understand?

OpenStudy (zzr0ck3r):

Use comparison to see it converges then do what i did to get the sum

OpenStudy (anonymous):

Thx

OpenStudy (zzr0ck3r):

@Moyo30 understand?

OpenStudy (zzr0ck3r):

Np

OpenStudy (anonymous):

comparison test using which function to compare?

OpenStudy (zzr0ck3r):

1/n^2

OpenStudy (anonymous):

thx, i have to go away. will look @ it when i come back.